# MambaCSR: Dual-Interleaved Scanning for Compressed Image Super-Resolution With SSMs

### v1

@@ A.2 Limitations
In this work, we introduce MambaCSR and propose dual-interleaved, position-aligned cross-scale scanning methods. While our approach demonstrates state-of-the-art performance on JPEG-compressed image super-resolution tasks, the generalization of MambaCSR to other compression standards remains unexplored. The CSR field encompasses a wide range of compression methods beyond JPEG, including traditional codecs such as VVC [3] and HEVC [43, 26], as well as newer, learning-based codecs [52, 17] like HIFIC [36]. The applicability of MambaCSR across these diverse compression techniques is still an open question. Future work will aim to evaluate MambaCSR’s performance on various compression standards and explore universal restoration methods. Additionally, the current training speed of MambaCSR for low-level tasks is relatively slow. Addressing this limitation will involve investigating more efficient training strategies and network designs to enhance MambaCSR’s training efficiency in future research.
  Datasets
Cost Set5 Set14 Manga109 Urban100 DIV2K Test
Methods Params FLops PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
SwinIR 8.72 37.17 26.51 0.7622 25.03 0.6393 24.89 0.7809 23.04 0.6462 26.72 0.7182
Swin2SR 11.67 52.96 26.56 0.7631 25.05 0.6394 24.92 0.7810 23.07 0.6471 26.73 0.7187
HST 16.48 54.95 26.53 0.7625 25.02 0.6395 24.91 0.7808 23.06 0.6467 26.73 0.7184
HAT 20.81 102.4 26.57 0.7624 25.04 0.6397 24.93 0.7809 23.12 0.6491 26.74 0.7196
MambaIR 16.70 82.30 26.61 0.7640 25.11 0.6400 25.01 0.7830 23.13 0.6530 26.76 0.7201
MambaCSR 19.18 80.46 26.67 0.7680 25.13 0.6420 25.12 0.7880 23.24 0.6580 26.79 0.7204
@@
Methods Params FLops PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
SwinIR 8.72 37.17 27.48 0.7849 25.61 0.6596 25.79 0.8054 23.59 0.6725 27.28 0.7355
Swin2SR 11.67 52.96 27.51 0.7846 25.61 0.6594 25.81 0.8057 23.63 0.6731 27.30 0.7359
HST 16.48 54.95 27.49 0.7851 25.60 0.6595 25.8 0.8058 23.61 0.6765 27.31 0.7361
HAT 20.81 102.4 27.53 0.7856 25.61 0.6608 25.84 0.8077 23.67 0.6778 27.32 0.7369
MambaIR 16.70 82.30 27.56 0.7860 25.64 0.6600 25.87 0.8080 23.72 0.7900 27.34 0.7375
MambaCSR 19.18 80.46 27.63 0.7880 25.70 0.6630 26.01 0.8120 23.82 0.6825 27.39 0.7398
[…]
@@ A.3 More Quantitative Results
In this section, we present additional quantitative results that extend beyond the main text, specifically focusing on settings with quality factors of 20 and 30. We compare our approach against four transformer-based models: SwinIR [28], Swin2SR [8], HST [23], HAT [5], as well as one Mamba-based model, MambaIR [15]. The evaluation is conducted on five compressed benchmarks: Set5 [2], Set14 [63], Manga109 [34], Urban100 [18], and the DIV2K test set [1]. All benchmarks are compressed using OpenCV library functions. As shown in Table 5 and Table 6, our model achieves optimal performance across all benchmarks, notably outperforming MambaIR by an average of approximately 0.11-0.12dB on the Manga109 and Urban100 datasets. Furthermore, our computational complexity remains comparable to that of MambaIR, which further demonstrates the efficiency and effectiveness of our approach.
A.4 More Visual Results
In this section, we present additional visual comparisons for QF=20 and QF=30, as illustrated in Fig.5. From the first image with a quality factor of 20, it is clear that our MambaCSR model excels at removing compression artifacts, providing a more orderly restoration of textures and structural details compared to other models. At a quality factor of 30, our model demonstrates its ability to recover finer-grained textures. In contrast, while MambaIR produces smoother overall images, it lacks detailed local structural information, underscoring its limitations in capturing localized features. We attribute these improvements to our proposed hierarchical-interleaved scanning and position-aligned cross-scale scanning methods, which offer superior contextual information modeling relative to other models.
Figure 5: Visual Comparisons tested on Urban100 [18] at x4 scale of QF=[20,30].

### v2

@@ A.2 Limitations
In this work, we introduce MambaCSR and propose dual-interleaved, position-aligned cross-scale scanning methods. While our approach demonstrates state-of-the-art performance on JPEG, HEVC and VVC compressed image super-resolution tasks, the generalization of MambaCSR to other compression standards remains unexplored. The CSR field encompasses a wide range of compression methods beyond JPEG, VVC [4] and HEVC [48] including traditional codecs such as WebP [16] and AVC [55], as well as newer, learning-based codecs [58, 23] like HIFIC [41]. The applicability of MambaCSR across these diverse compression techniques is still an open question. Future work will aim to evaluate MambaCSR’s performance on various compression standards and explore universal restoration methods. Additionally, the current training speed of MambaCSR for low-level tasks is relatively slow. Addressing this limitation will involve investigating more efficient training strategies and network designs to enhance MambaCSR’s training efficiency in future research.
@@
Codec Methods Params Flops PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
SwinIR [33] 8.72 37.17 29.06 0.822 26.84 0.709 28.12 0.860 24.83 0.730 28.64 0.795
Swin2SR [9] 11.67 52.96 29.04 0.822 26.82 0.706 28.13 0.858 24.81 0.727 28.63 0.793
HST [29] 16.48 54.96 29.07 0.821 26.79 0.708 28.11 0.857 24.79 0.726 28.65 0.795
HAT [6] 20.81 102.4 29.09 0.824 26.87 0.711 28.19 0.862 24.97 0.737 28.67 0.796
MambaIR [21] 16.70 82.3 29.11 0.823 26.87 0.710 28.22 0.861 24.96 0.734 28.68 0.797
VVC (QP=32) MambaCSR 19.69 76.03 29.15 0.827 26.93 0.714 28.31 0.864 25.08 0.741 28.73 0.799
SwinIR [33] 8.72 37.17 27.21 0.776 23.53 0.658 26.27 0.818 23.68 0.679 26.99 0.756
Swin2SR [9] 11.67 52.96 27.23 0.777 25.51 0.654 26.25 0.818 23.67 0.679 26.96 0.756
HST [29] 16.48 54.96 27.20 0.776 25.52 0.657 26.25 0.817 23.67 0.678 27.00 0.755
HAT [6] 20.81 102.4 27.24 0.779 25.55 0.658 26.30 0.819 23.74 0.682 27.01 0.757
MambaIR [21] 16.70 82.3 27.26 0.779 25.57 0.659 26.32 0.821 23.77 0.684 27.03 0.758
VVC (QP=37) MambaCSR 19.69 76.03 27.30 0.783 25.62 0.661 26.43 0.824 23.85 0.686 27.06 0.761
SwinIR [33] 8.72 37.17 25.12 0.721 23.93 0.603 24.08 0.762 22.27 0.615 24.73 0.657
Swin2SR [9] 11.67 52.96 25.11 0.720 23.91 0.602 24.06 0.761 22.25 0.612 24.72 0.655
HST [29] 16.48 54.96 25.14 0.722 23.95 0.603 24.08 0.760 22.28 0.616 24.74 0.657
HAT [6] 20.81 102.4 25.14 0.722 23.95 0.604 24.12 0.763 22.31 0.619 24.74 0.658
MambaIR [21] 16.70 82.3 25.16 0.724 23.99 0.605 24.20 0.765 22.25 0.622 24.76 0.659
VVC (QP=42) MambaCSR 19.69 76.03 25.18 0.725 24.03 0.606 24.26 0.767 22.42 0.621 24.77 0.661
@@ A.3 More Quantitative Results
In this section, we present additional quantitative results that complement the main text, focusing specifically on the VVC codec under QP settings of [32, 37, 42]. Our approach is compared against four transformer-based models: SwinIR [33], Swin2SR [9], HST [29], and HAT [6], as well as the Mamba-based model, MambaIR [21]. We adopt DF2K as our primary training dataset [2]. The evaluation is performed on five compressed benchmarks: Set5 [3], Set14 [69], Manga109 [39], Urban100 [24], and the DIV2K test set [2], with all datasets compressed using VTM software.
As shown in Table 5, our model achieves state-of-the-art performance across all benchmarks, significantly surpassing MambaIR by an average of approximately 0.11–0.12 dB on the Manga109 and Urban100 datasets. Moreover, our approach reduces computational complexity by 6 GFLOPS compared to MambaIR, underscoring both the efficiency and effectiveness of our proposed method.
A.4 More Visual Comparisons
In this section, we provide additional visual comparisons for the HEVC and VVC codec across three distortion levels: QP values from [32, 37, 42]. As depicted in Fig. 5 and Fig. 6, it is evident that our MambaCSR model demonstrates a significant advantage in mitigating compression artifacts, delivering a more structured restoration of textures and structural details compared to other approaches. In contrast, MambaIR falls short in preserving detailed local structural information, highlighting its limitations in capturing fine-grained features. These advancements can be attributed to our proposed hierarchical-interleaved scanning and position-aligned cross-scale scanning strategies, which enable more superior contextual information modeling than other existing models.
Figure 5: Visual Comparisons tested on Urban100 [24] for HEVC codec [48] at x4 scale of QP=[32, 37, 42].
Figure 6: Visual Comparisons tested on Urban100 [24] for VVC codec [4] at x4 scale of QP=[32, 37, 42].